9. Solve the ILLP given below by graphical method: (10)

Maximum 95 1 100 2 Z = x + x

Subject to the Constraints

5x1 + 2x2 ≤ 20

x1 ≥ 3

5 x2 ≤

1 2

x x are non-negative integers.


1
Expert's answer
2021-05-07T09:48:01-0400

Solution:

Maximize "Z=95x_1+100x_2"

subject to the constraints:

"5x_1+2x_2≤ 20 ...(i)
\\x_1 ≥ 3...(ii)
\\x_2 ≤ 5...(iii)"

"x_1 , x_2\ge0"

Consider (i), (ii), (iii) equations and plotting their graph, we get,



Clearly, corner points are "(3,0),(4,0),(3,2.5)".

At "(3,0),Z=95(3)+100(0)=285"

At "(4,0),Z=95(4)+100(0)=380"

At "(3,2.5),Z=95(3)+100(2.5)=535"

Hence, maximum value is 535 at "(3,2.5)".